40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Nov 12 18:08:35 2014, Jenne, Update, LSC, RIN in transmission a problem? RIN_TRX_TRY_MCtrans_12Nov2014_zip.xml.gzRIN_TRX_TRY_MCtrans_12Nov2014.pdfRINcontribution_12Nov2014_linearXscale.pngRINcontribution_12NOv2014_logXscale.png
    Reply  Thu Nov 13 01:00:37 2014, rana, Update, LSC, RIN in transmission a problem? 
    Reply  Thu Nov 13 04:28:28 2014, Jenne, Update, LSC, RIN in transmission a problem? zippyzip.zipRIN_TRX_TRY_IPPOS_NoLens_12Nov2014.pdf
       Reply  Thu Nov 13 23:42:01 2014, Jenne, Update, LSC, RIN vs. Seismic Seismic_TRXTRYandMC_13Nov2014.pdf
          Reply  Fri Nov 14 17:08:17 2014, Jenne, Update, LSC, RIN vs. Seismic 
       Reply  Fri Nov 14 20:31:13 2014, ericq, Update, LSC, RIN in transmission a problem? 
Message ID: 10709     Entry time: Thu Nov 13 04:28:28 2014     In reply to: 10703     Reply to this: 10712   10720
Author: Jenne 
Type: Update 
Category: LSC 
Subject: RIN in transmission a problem? 

[Jenne, Rana, Koji]

We did some thinking on what could be causing the excess RIN that we see in the arm transmissions but not in the MC transmission.  Unfortunately, I don't think we have anything conclusive yet. 

We thought about:

  • Test mass oplevs
  • Input tip tilt jitter
  • MC motion

Oplevs

As Rana reported in elog 10708, we tuned the oplev servos for ITMX, ETMX, ITMY and ETMY.  They all now look like the new SRM oplevs that Rana described in elog 10694.  However, when we re-looked at the RIN after the oplev tuning, we did not see a noticeable change.  So, fixing up the oplevs didn't fix up the RIN.

Side notes:

  • Several optics had low gains for suspos, which were increased to give Qs of ~5ish.
  • When we gave ITMX a 500 count step in pitch (the same size used for all other optics in both pit and yaw), it didn't come back afterward.  This is a little disconcerting.  Rana had to move the alignment slider to get it back so that we had MICH fringing at the AS port again.

Input Tip Tilts

Koji did some work, reported in elog 10706, on how much the jitter of the input pointing tip tilts should affect us.  We don't think that they are moving enough to be the cause of the excess RIN that we see.


Mode Cleaner Motion

We see some coherence between MC2 suspit and TRX/TRY, so we suspect that the MC's motion could be causing problems. 

I looked at the WFS vs. TRX & TRY, and saw significant coherence at the 3 Hz stack resonance.  I think it's clear that the WFS can help suppress this motion more.  The WFS noise level was too bad to see any other coherence at other frequencies. (Side note:  We should consider increasing the analog WFS signal.  As Rana mentioned back in 2008 in elog 454, the signal is super small.  Increasing the analog gain could allow us to suppress motion at more frequencies, although it would be a bit of a pain.)

To try and get some more signal, I routed the IPPOS beam over to the PRM oplev temporarily.  The idea was to be able to look at the IPPOS port, but with a fast channel.  I turned off the BS/PRM HeNe, and removed the last steering mirror before the QPD.  I put in 2 Y1 steering mirrors to get the IPPOS beam across the table and pointing at IPPOS.  I took my measurements 3 times, with different beam sizes on the QPD, to try to image different gouy phases.  I used absorptive ND filters (0.6 + 0.1) to get the light level on the PD such that I had about 10,000 counts per quadrant, where 16,000 counts seemed to be the saturation point. I also reset the dark offsets of the QPD quadrants, although they were so small that I don't think it did much.  I also took out the optical lever calibration from counts to microradians, since that number isn't meaningful for what I was doing.  So, the IPPOS signals (using the PRM oplev channels) are in raw ADC counts.  The first measurement had no lens, and the beam was probably at least half the size of the QPD.  The second measurement had a lens, and the beam on the QPD was about half the original size.  The third measurement had the lens closer to the QPD, so that the beam was about 1mm on the diode.  In none of these cases do I see any significant coherence with the TRX/TRY RIN signals, except at 3 Hz. After my measurements I put the oplev back including all of the digital settings, although for now I just left the IPPOS beam dumped on a razor dump, since it wasn't being used anyway.  I need to realign IPPOS when it's not the middle of the night.

Some thoughts that we have:

  • Maybe it's time to resurrect seismic feedforward on MC length, to suppress some of this 3 Hz motion?
  • Maybe we should be using the MC_L path to offload some of the frequency feedback, so that we're not pushing on MC2 so hard (because if we have bad F2P coupling, this creates beam motion)

I have plots for each of my IPPOS vs. TRX/TRY measurements.  The data is attached.  For each, I looked at the transfer function between IPPOS (using the SUS-PRM_OPLEV channels) and TRX/TRY to get the 'calibration' between input beam motion and transmission RIN.  In all cases, at 3 Hz the coefficient was about 1, so in the power spectra on the right side, there is no calibration applied to the IPPOS signals. 

IPPOS vs. Transmission RIN, no lens, big beam on QPD:

RIN_TRX_TRY_IPPOS_NoLens_12Nov2014.pdf

(Just kidding about the other 2 plots - the elog can't handle it.  They're in the zippyzip file though, and I don't think they look qualitatively different from the no-lens case).

 

Attachment 1: zippyzip.zip  31.401 MB
ELOG V3.1.3-